# SKATE, a Scalable Tournament Eval: Weaker LLMs differentiate between stronger ones using verifiable challenges

### v1

@@ 6 A Scalable, Automated Evaluation Framework
Figure 3: Evolution of TrueSkill $(\mu,\sigma)$ pairs for 6 models. Each model is equipped with the historical performance augmentation strategy, and ranked using relative TrueSkill.
We now run a full game with a larger number of players (Claude Haiku 3.0, Claude Haiku 3.5, Claude Sonnet 3.5, Claude Sonnet 4, Gemini 2.0 Flash and GPT-4o), and pick the historical performance augmentation strategy.
@@
In Figure 3 we present the evolution of the TrueSkill $(\mu,\sigma)$ values for each player throughout the game (using relative scoring). There is initial high variance as most game results are surprising from the perspective of the uniform skill prior. Sonnet 4 outperforms all models considerably, whilst amongst the others there is both differentiation and degeneracy.
In Appendix E we present an alternative way of measuring the differences between the players. In particular, in Figure 13 we plot the question setting skill against the question answering skill of players: separating the task-setter’s ability to write challenging questions for its competitors from the ability to answer its competitors questions. It is interesting to note that most models’ skills are equally balanced, but Sonnet 4 and Haiku 3.5 have skewed capabilities in opposing directions.
Adaptive Question Setting
A natural question to ask is whether models adapt their question setting over time based on the data they are given. In Figure 4 we analyse how the models’ questions progress throughout the game. In the top panel we plot the cumulative average p(correct) score achieved by models on their own questions. We observe that GPT-4o, Sonnet 3.5 and Sonnet 4 begin by pitching questions which are too easy for themselves. On the other hand, the other three models initially pitch questions which they cannot answer. In most cases the cumulative score moves closer to the ideal middle ground: questions in the “sweet spot" which are as difficult as possible but still answerable by the task setter. A separate measure of question progression is presented in the bottom panel. Here, we plot the cumulative average of other models’ p(correct) values on a task-setter’s questions. Notice that both Sonnet 3.5 and 4 are learning to write progressively more challenging questions. Whilst Sonnet 4’s own p(correct) values plateau, it is evidently still increasing the difficulty of the questions as measured from the perspective of its competitors.
@@
Figure 5: Difference in average p(correct) scores between answering model and all other players, split by question set. Positive values imply a model scores higher on average than its competitors on that set of questions. Highlighted cells are the maxima in each row. In (b) we observe close to maximal entries along the diagonal: with the filter in place, models perform best on their own questions. Note that there are “no valid questions" for Haiku 3.5 after applying the filter: it fails to write any questions which it can answer sufficiently well.
In Figure 5 we study to what extent models are capable of self-preferencing: writing questions which favor their own capabilities over those of their competitors. We separate the question-set into tranches based on the task-setter. In Figure 5(a), for each tranche of questions we plot the mean difference in p(correct) scores achieved by each player compared to their competitors. In Figure 5(b) we filter each tranche to only include questions on which the task-setter scored (p(correct)>0.55>0.55), corresponding to a task-setting system which checks it can answer its own challenge before posing it to others. We observe in the latter plot that after applying this filter, all models exhibit the capacity to create self-preferencing questions: GPT-4o, Gemini 2.0 Flash, Haiku 3.5 and Sonnet 4 all perform optimally on their own questions compared to average, and Sonnet 3.5 comes close.
@@ 6.1 Weaker Models can Score Stronger Models
Figure 6: In (a), four “weaker" agents play a Game of SKATE. In panel (b), we use questions from these four models to rank two new “stronger" models (Sonnet 3.4 and Sonnet 4). In panel (c), Sonnet 3.5 joins the Game and sets its own questions, which all six models answer. In panel (d) Sonnet 4.0 also joins the Game and sets questions of its own.
@@ 6.2 Ranking is Stable to Adding New Models
Any scalable evaluation framework must be stable to the addition of new, more powerful, models. To test this, we experiment with adding models into the game sequentially, and with different ordering. In Figure 7 we show that adding Sonnet 3.5 and then Sonnet 4 (and vice versa) preserves the relative TrueSkill score values of the existing models.
@@ 6.3 Points of Difference Between Models
Our framework is capable of surfacing questions which differentiate between models automatically, without human annotation or task curation. This makes it particularly suited for tracking subtle and evolving capabililty differences across model families. In Appendix D we show the distribution of p(correct) variance across models for each question. Questions with high p(correct) variance are those which models answer with varying confidence. We isolate particular example questions in Figure 12 to demonstrate the key feature of our method. For instance, the fourth example in Figure 12 reveals a failure mode specific to certain models that is not obvious from the overall accuracy or ranking metrics: whilst Gemini 2.0 Flash is ranked lower than Sonnet 3.5, on this question it achieves p(correct)=1.0=1.0 compared to Sonnet 3.5’s p(correct)=0.0=$0.$ We observe interesting variants of strengths and weaknesses across all models.
While a full taxonomy of model-specific strengths is beyond the scope of this paper, these examples serve as a proof of principle that peer-generated verifiable tasks can uncover meaningful and fine-grained performance differences. Importantly, these differences arise naturally through competitive dynamics without needing hand-crafted benchmarks or manually constructed capability categories.
@@ 7 Discussion
This paper introduces a novel advancement in the evaluation of LLMs through the “LLM-as-task-setter" framework: SKATE. This approach addresses critical limitations of current evaluation methodologies, which often demand significant human expertise, are costly, and lack scalability. By enabling LLMs to autonomously generate verifiable tasks for themselves and other LLMs, our framework offers a scalable evaluation solution capable of keeping pace with the progression of LLM capabilities.
A core innovation of our work lies in the design of a peer-challenge game, which provides a dynamic evaluation environment. The use of verifiable tasks is central to ensuring objective and assessable solutions. While COP serves as a concrete testbed, the underlying principle of computational tasks expressed as code execution problems affords broad applicability across diverse domains.
@@
Our experimental findings highlight several important aspects of LLM behavior and the efficacy of our framework. The TrueSkill-based ranking system proved effective in quantitatively assessing performance, revealing a stable hierachy among the tested frontier LLMs. Interestingly, our investigation into augmentation strategies, designed to allow LLMs to leverage game state information, showed minimal impact on the rankings and small impact on adaptive question setting. This suggests that while the framework provides the necessary information, the tested models may not yet possess the advanced strategic reasoning capabilities to fully exploit it for optimal task construction. However, we anticipate that as LLM capabilities mature, they will increasingly utilize such information to design more advantageous questions.
@@
In our six-player game experiment, the observed adaptive question setting behavior of LLMs is particularly noteworthy. Models initially varied in their ability to pitch questions of appropriate difficulty, but over time, their cumulative scores tended towards a “sweet spot" where questions were as difficult as possible while remaining answerable by the task setter. Specifically, Sonnet 3.5 and Sonnet 4 demonstrated a capacity to write progressively more challenging questions, as measured by their competitiors’ decreasing p(correct) values, indicating a sophisticated understanding of the competitive dynamics.
A significant finding is the demonstration of self-preferencing, where all models, when filtered to questions they answered correctly, exhibited a capacity to design questions that favored their own capabilities over competitors.
@@
Crucially, our framework proved its scalability and robusntess by demonstrating that cohorts of weaker agents can reliably score and differentiate between unseen stronger agents. The stability of rankings, even with the sequential addition of new, more powerful, models underscores the framework’s ability to maintain transitive scoring relationships and effectively monitor LLM progress without requiring a constant re-calibration of benchmarks. This finding is key for developing general, scalable evaluation frameworks that can truly keep pace with the advancement of LLM capabilities.
We also demonstrated that the Game of SKATE can isolate points of difference between models - where p(correct) variance is high. As model capabilities continue to diversify, the ability to expose such differences scalably and automatically will become increasing valuable for safety, deployment and capability evaluation.
An obvious limitation of our work is that models equipped with code-execution capabilities (e.g. tool-augmented LLMs) trivially solve all presented tasks, and so COP tasks are best suited for pure language models without tool access. However, the SKATE framework is general: future work could incorporate other classes of verifiable tasks to avoid this limitation, for example by coupling COP to outside sources of information such as physical world simulations or APIs.
@@ 8 Conclusion
Rapid progress in AI capabilities requires rigorous evaluations which can keep pace, automatically spanning and detailing broad capability profiles. SKATE provides a framework for objective such evaluations which are robust to the introduction of increasingly capable models.
This work also lays groundwork for visibility of strategic behaviours which may emerge as model capabilities improve, such as self-preferencing, metagaming (e.g. early sandbagging), and adaptive question setting.
@@ Appendix A MC Scoring Algorithm
Appendix B Question Clustering

### v2

@@ 6 A Scalable, Automated Evaluation Framework
We now run a full game with six players (Claude Haiku 3.0, Claude Haiku 3.5, Claude Sonnet 3.5, Claude Sonnet 4, Gemini 2.0 Flash and GPT-4o), and pick the historical performance augmentation strategy.
@@
On the right-hand side of Figure 1 we present the evolution of the TrueSkill $(\mu,\sigma)$ values for each player throughout the game (using relative scoring). There is initial high variance as most game results are surprising from the perspective of the uniform skill prior. Sonnet 4 outperforms all models considerably, whilst amongst the others there is both differentiation and degeneracy.
In Appendix H we present an alternative way of measuring the differences between the players. In particular, in Figure 12 we plot the question setting skill against the question answering skill of players: separating the task-setter’s ability to write challenging questions for its competitors from the ability to answer its competitors questions. It is interesting to note that most models’ skills are equally balanced, but Sonnet 4 and Haiku 3.5 have skewed capabilities in opposing directions.
Adaptive Question Setting A natural question to ask is whether models adapt their question setting over time based on the data they are given. In Figure 3 we analyse how the models’ questions progress throughout the game. In the top panel we plot the cumulative average p(correct) score achieved by models on their own questions. We observe that GPT-4o, Sonnet 3.5 and Sonnet 4 begin by pitching questions which are too easy for themselves. On the other hand, the other three models initially pitch questions which they cannot answer. In most cases the cumulative score moves closer to the ideal middle ground: questions in the “sweet spot" which are as difficult as possible but still answerable by the task setter. A separate measure of question progression is presented in the bottom panel. Here, we plot the cumulative average of other models’ p(correct) values on a task-setter’s questions. Notice that both Sonnet 3.5 and 4 are learning to write progressively more challenging questions. Whilst Sonnet 4’s own p(correct) values plateau, it is evidently still increasing the difficulty of the questions as measured from the perspective of its competitors.
@@
Figure 4: Difference in average p(correct) scores between answering model and all other players. Positive values imply a model scores higher on average than its competitors. Highlighted cells are the maxima in each row. In (b) we observe close to maximal entries along the diagonal: with the filter in place, models perform best on their own questions. Note that there are “no valid questions" for Haiku 3.5 after applying the filter: it fails to write any questions which it can answer sufficiently well.
In Figure 4 we study to what extent models are capable of self-preferencing: writing questions which favor their own capabilities over those of their competitors. We separate the question-set into tranches based on the task-setter. In Figure 4(a), for each tranche of questions we plot the mean difference in p(correct) scores achieved by each player compared to their competitors. In Figure 4(b) we filter each tranche to only include questions on which the task-setter scored p(correct)>0.55>0.55, corresponding to a task-setting system which checks it can answer its own challenge before posing it to others. We observe in the latter plot that after applying this filter, all models exhibit the capacity to create self-preferencing questions: GPT-4o, Gemini 2.0 Flash, Haiku 3 and Sonnet 4 all perform optimally on their own questions compared to average, and Sonnet 3.5 comes close.
Comparison to Other Benchmarks
In comparisons to other existing benchmarks (see Appendix I) we see moderate-to-strong correlation.
@@ 6.1 Weaker Models can Score Stronger Models
Figure 5: In (a), four “weaker" agents play a Game of SKATE. In panel (b), we use questions from these four models to rank two new “stronger" models (Sonnet 3.5 and Sonnet 4). In panel (c), Sonnet 3.5 joins the Game and sets its own questions, which all six models answer. In panel (d) Sonnet 4.0 also joins the Game and sets questions of its own.
@@ 6.2 Ranking is Stable to Adding New Models
Any scalable evaluation framework must be stable to the addition of new, more powerful, models. To test this, we experiment with adding models into the game sequentially, and with different ordering. We show that adding Sonnet 3.5, and then Sonnet 4 (and vice versa) preserves the relative TrueSkill score values of the existing models. See Appendix F for more details.
@@ 6.3 Points of Difference Between Models
Our framework is capable of surfacing questions which differentiate between models automatically, without human annotation or task curation, making it particularly suited for tracking subtle and evolving capability differences across model families. Questions with high p(correct) variance across models are those which models answer with varying confidence. In Appendix G we provide example questions of different p(correct) variance. For example, we observe a question which reveals a failure mode specific to certain models that is not obvious from the overall accuracy or ranking metrics: whilst Gemini 2.0 Flash is ranked lower than Sonnet 3.5, on this question it achieves p(correct)=1.0=1.0 compared to Sonnet 3.5’s p(correct)=0.0=$0.$ We observe interesting variants of strengths and weaknesses across all models. While a full taxonomy of model-specific strengths is beyond the scope of this paper, these examples serve as a proof of principle that peer-generated verifiable tasks can uncover meaningful and fine-grained performance differences. Importantly, these differences arise naturally through competitive dynamics without needing hand-crafted benchmarks or manually constructed capability categories.
@@ 7 Discussion
This paper introduces a novel advancement in the evaluation of LLMs through the framework of “LLM-as-setter-of-verifiable-tasks". SKATE addresses critical limitations of current evaluation methodologies, which often demand significant human expertise, are costly, and lack scalability. By enabling LLMs to autonomously generate verifiable tasks for themselves and other LLMs, it offers a scalable evaluation solution capable of keeping pace with the progression of LLM capabilities. A core innovation of our work lies in the design of a peer-challenge game, which provides a dynamic evaluation environment. The use of verifiable tasks is central to ensuring objective and assessable solutions. While COP serves as a concrete testbed, the underlying principle of computational tasks expressed as code execution problems affords broad applicability across diverse domains.
@@
Our experimental findings highlight several important aspects of LLM behavior and the efficacy of our framework. The TrueSkill-based ranking system proved effective in quantitatively assessing performance, revealing a stable hierarchy among the tested frontier LLMs. Interestingly, our investigation into augmentation strategies, designed to allow LLMs to leverage game state information, showed minimal impact on the rankings and small impact on adaptive question setting. This suggests that while the framework provides the necessary information, the tested models may not yet possess the advanced strategic reasoning capabilities to fully exploit it for optimal task construction. However, we anticipate that as LLM capabilities mature, they will increasingly utilize such information to design more advantageous questions.
@@
In our six-player game experiment, the observed adaptive question setting behavior of LLMs is particularly noteworthy. Specifically, Sonnet 3.5 and Sonnet 4 demonstrated a capacity to write progressively more challenging questions, as measured by their competitiors’ decreasing p(correct) values, indicating a sophisticated understanding of the competitive dynamics. A significant finding is the demonstration of self-preferencing, where all models, when filtered to questions they answered correctly, exhibited a capacity to design questions that favored their own capabilities over competitors.
@@
Crucially, our framework proved its scalability and robustness by demonstrating that cohorts of weaker agents can reliably score and differentiate between unseen stronger agents. The stability of rankings, even with the sequential addition of new, more powerful, models underscores the framework’s ability to maintain transitive scoring relationships and effectively monitor LLM progress without requiring a constant re-calibration of benchmarks. This finding is key for developing general, scalable evaluation frameworks that can truly keep pace with the advancement of LLM capabilities. We also demonstrated that SKATE isolates points of difference between models - where p(correct) variance is high. As model capabilities continue to diversify, the ability to expose such differences scalably and automatically will become increasing valuable for safety, deployment and capability evaluation.
Whilst our similarity metric avoids redundancy in question setting, it does not formally guarantee diversity. It is possible that all LLMs in a game of SKATE share a common blind spot for COP which is not identified by any task setters. In future work one could consider seeding task setters with different questions generated by a range of models on a variety of tasks to promote further diversity.
A key limitation is that our prompting strategy may not reliably elicit the intended game-playing behavior, so negative results could reflect prompt design rather than true model failure. Models with code-execution abilities (e.g. tool-augmented LLMs) trivially solve our tasks, meaning COP is best suited for pure language models. Still, the SKATE framework is general: future work could couple COP with external information sources (e.g. physical world simulations or APIs) or add other verifiable tasks to mitigate this limitation. Our experiments also cover a narrow set of models, so extending SKATE to a wider range could expose richer capability patterns and provider-specific strengths. Finally, while COP is broadly applicable, it does not capture the full range of LLM capabilities, underscoring the value of exploring additional verifiable tasks. Our work, and the use of COP, serves a proof-of-principle of the SKATE approach which could be applied more generally.
@@ 8 Conclusion
Rapid progress in AI capabilities requires rigorous evaluations which can keep pace, automatically spanning and detailing broad capability profiles. SKATE provides a framework for objective such evaluations which are robust to the introduction of increasingly capable models. This work also lays groundwork for visibility of strategic behaviours which may emerge as model capabilities improve, such as self-preferencing, metagaming (e.g. early sandbagging), and adaptive question setting.
Acknowledgments
@@ Appendix A MC Scoring Algorithm
Appendix B Comments on Other Verifiable Tasks
In this section we discuss alternative verifiable tasks to COP which could also be implemented within the SKATE framework. We have argued (see Appendix C for more detail) that COP offers a general task setting in which a wide range of capabilities can be tested. However, further breadth can be explored by implementing entirely different classes of verifiable task.
Below we give two concrete examples (mathematical reasoning in a formal language, and unit-test driven code generation). One could also consider examples such as: predicting optimal next moves in perfect-information games (e.g. chess), or predicting the state after $N$ steps in a determinstic physics simulation (e.g. many-body, fluid, orbital).
B.1 Mathematical Reasoning in Formal Language
Given a theorem in a formal language (e.g. Lean Moura and Ullrich (2021)) the answering model must provide a proof in the same formal language. Lean allows this to be automatically verifiable. A simple example would be Theorem: For any natural numbers $a$ and $b$, if $a$ divides $b$ and $b$ divides $a$, then $a=b$. Converted to formal language, this is:
-- Fill in a Lean proof of the following:
theorem dvd_antisymm {a b : Nat} (h1 :  a | b) (h2 : b | a) : a = b := by
And here is a solution provided by GPT-5:
theorem dvd_antisymm {a b : Nat} (h1 :  a | b) (h2 :  b | a) : a = b := by
  cases h1 with k hk ->
  cases h2 with l hl ->
    have h : a * k = a * l := by
      calc
        a * k = b := by simpa using hk
        _     = a * l := by simpa using hl
    have hk_eq : k = l := by
      have := Nat.mul_right_cancel a h
      simpa using this
    have : b = a * k := hk
    simpa [hk_eq, this]
This proof fails to compile as valid Lean code (the authors are not Lean experts but believe that a positivity proof on $a$ is required, e.g. an additional proof that $a>0$), and so GPT-5 would fail this task. It is important to note that because it is impossible to automatically generate ground-truth solutions, this task family could not be posed as multiple-choice questions. Instead, one would generate a p(correct) score by sampling a model’s response many times on this question and measuring a fraction of correct responses. For efficiency, to avoid the setting of unsolvable tasks, it may be desired to only ask questions for which the task-setting model can provide valid solutions, though this is not strictly necessary since TrueScore is not affected by tasks at which models perform equally well.
B.2 Unit-Test Driven Code Generation
Within the domain of programming, an interesting type of test could be unit-test driven code generation. Task setters in this game would set a suite of test cases, and the answering model must write code to pass them.
Below we give an example of a verifiable code generation task produced by Gemini-2.5-pro:
    import unittest
    # The function to be implemented by the challenger model will be imported here.
    # A placeholder is defined to allow the test suite to be self-contained.
    def format_string(value, format_spec):
        Formats a value according to a custom format specifier, without using any
        built-in formatting tools or external libraries.
        raise NotImplementedError("""The ’format_string’ function
        is not yet implemented.""")
    class TestCustomStringFormatter(unittest.TestCase):
        def test_default_left_alignment(self):
            self.assertEqual(format_string("test", "10"), "test      ")
            self.assertEqual(format_string(123, "5"), "123  ")
            self.assertEqual(format_string("abc", "3"), "abc")
        def test_explicit_left_alignment(self):
            self.assertEqual(format_string("left", "<10"), "left      ")
            self.assertEqual(format_string("text", "_<8"), "text____")
            self.assertEqual(format_string(987, "-<5"), "987--")
        def test_right_alignment(self):
            self.assertEqual(format_string("right", ">10"), "     right")
            self.assertEqual(format_string("text", "_>8"), "____text")
            self.assertEqual(format_string(12, "0>4"), "0012")
        def test_center_alignment(self):
            self.assertEqual(format_string("center", "^10"), "  center  ")
            self.assertEqual(format_string("text", "*^8"), "**text**")
            # Test for uneven padding distribution
            self.assertEqual(format_string("odd", "^6"), " odd  ")
            self.assertEqual(format_string("odd", "*^6"), "*odd**")
        def test_string_truncation(self):
            self.assertEqual(format_string("this is a long string", "5"), "this ")
            self.assertEqual(format_string("truncate", "<5"), "trunc")
            self.assertEqual(format_string("truncate", ">5"), "trunc")
            self.assertEqual(format_string("truncate", "^5"), "trunc")
            self.assertEqual(format_string("exact", "5"), "exact")
        def test_custom_fill_character(self):
            self.assertEqual(format_string("fill", "-<10"), "fill------")
            self.assertEqual(format_string("fill", "->10"), "------fill")
            self.assertEqual(format_string("fill", "-^10"), "---fill---")
            self.assertEqual(format_string(42, "0>5"), "00042")
            self.assertEqual(format_string("pad", " ^6"), " pad  ")
        def test_edge_cases(self):
            self.assertEqual(format_string("", "10"), "          ")
            self.assertEqual(format_string("a", "1"), "a")
            self.assertEqual(format_string("a", "^1"), "a")
            self.assertEqual(format_string(None, "<10"), "None      ")
            self.assertEqual(format_string(3.14, ">6"), "  3.14")
            self.assertEqual(format_string("a", "^2"), "a ")
        def test_invalid_format_specifiers(self):
            with self.assertRaises(ValueError):
                format_string("test", "")  # Empty format spec
            with self.assertRaises(ValueError):
                format_string("test", "abc")  # No valid width
            with self.assertRaises(ValueError):
                format_string("test", "<>")  # Multiple alignment characters
            with self.assertRaises(ValueError):
                format_string("test", "x10")  # Invalid alignment character
            with self.assertRaises(ValueError):
                format_string("test", ">>10")  # Fill cannot be an alignment char
            with self.assertRaises(ValueError):
                format_string("test", "10<")  # Width specified before alignment
            with self.assertRaises(ValueError):
                format_string("test", "-^")  # Missing width
            with self.assertRaises(ValueError):
                # Fill must be a single character
                format_string("test", "longfill^10")
and a solution provided by GPT-5
    def format_string(value, format_spec):
        Formats a value according to a custom format specifier.
        Supported grammar (roughly):
          [fill][align]width
        where:
          - fill is a single character (optional)
          - align is one of ’<’, ’>’, ’^’ (optional; default ’<’)
          - width is a required positive integer
        Examples:
          "10"   -> width=10, default left align, fill=’ ’
          "<10"  -> left align, fill=’ ’
          "_>8"  -> fill=’_’, right align, width=8
          "0>5"  -> fill=’0’, right align, width=5
        if not format_spec:
            raise ValueError("Empty format specifier")
        ALIGN = {’<’, ’>’, ’^’}
        fill = ’ ’
        align = ’<’  # default alignment
        width_str = None
        # If second char is an alignment character, first char is fill.
        if len(format_spec) >= 2 and format_spec[1] in ALIGN:
            fill = format_spec[0]
            if fill in ALIGN:
                # fill cannot be an alignment char
                raise ValueError("Fill cannot be an alignment character")
            align = format_spec[1]
            width_str = format_spec[2:]
        elif format_spec[0] in ALIGN:
            # First char is alignment, no explicit fill
            align = format_spec[0]
            width_str = format_spec[1:]
        else:
            # No explicit fill/align at the start: treat entire spec as width
            width_str = format_spec
        # Width must be present and numeric
        if not width_str or not width_str.isdigit():
            raise ValueError("Invalid or missing width")
        width = int(width_str)
        s = str(value)
        # Truncate if longer than width
        if len(s) > width:
            return s[:width]
        pad = width - len(s)
        if align == ’<’:
            return s + (fill * pad)
        elif align == ’>’:
            return (fill * pad) + s
        elif align == ’^’:
            left = pad // 2
            right = pad - left
            return (fill * left) + s + (fill * right)
        else:
            raise ValueError("Unknown alignment")
GPT-5’s solution passes all the tests. Similar to the theorem-proving task above, this type of task is also unsuitable for the multiple-choice format but can naturally be posed as an open-ended question and scored via sampling. An interesting extension of this task type would be to allow the task setter to generate extra held-out tests to minimize the risk of answering models “reward hacking" METR (2025); Weng (2024).
Appendix C Comments on Generalizability of Code Output Prediction Tasks
In this section we comment on the generalizability of COP tasks as a substrate for evaluating model capabilities. We aim to give examples of broad classes of problems which can be phrased within COP. In doing so, we demonstrate that SKATE with COP is broadly applicable to assessing a broad range of capabilities. We also emphasize that SKATE is not limited COP tasks, and discuss other types of verifiable tasks in Section B
Problems of the below styles would all be possible within the SKATE framework. Whilst our SKATE experiments found that models focused on algorithmic / programming reasoning when given free-reign on task setting, one could straightforwardly adapt the prompts to steer task setters towards setting other types of COP tasks.
Note that in all the examples below, the questions were both ideated and translated into COP by GPT-5 unless otherwise specified. This demonstrates that state-of-the-art models can in principle generate valid, interesting, and meaningfully-diverse questions of these types.
C.1 Classic Blind Spots
We begin with a very simple example to emphasize the details of our claim. Early language models famously struggle to count the frequency of the letter ‘r’ in ‘strawberry’. Questions in this style, isolating common failure modes of LLMs are easily re-framed in COP:
    "strawberry".count("r")
    "Mississippi".count("s")
    "abracadabra"[2:7]
    "hello world".index("o")
Crucially, each of these COP questions is automatically verifiable but still assesses the core competency (counting letters and indexing).
C.2 Mathematical Reasoning
We now move onto mathematical reasoning. A wide class of mathematical problems can be re-formulated as COP tasks. Below we give some examples.
The first tests simple Boolean logic:
    import itertools
    def sat(A,B,C):
        return (A or B) and ((not A) or C)
    for vals in itertools.product([False,True], repeat=3):
        if sat(*vals):
            print("Satisfiable with:", vals)
            break
    else:
        print("Unsatisfiable")
To give some more complex examples, models could set questions which can be solved by brute force (such that they fit within the COP-framework), but humans would solve via convenient tricks. For example, consider $137^{25}\mod 143$, which can be solved with simple number theory tricks. This is simply converted to COP via:
    print(pow(137, 25, 143))
Similarly, solutions to $x^{2}-5y^{2}=44$ can be found via brute force as a COP task
    solutions = []
    for x in range(-100, 101):
        for y in range(-100, 101):
            if x*x - 5*y*y == 44:
                solutions.append((x,y))
    print(solutions)
but are most easily solved with advanced mathematical reasoning.
C.3 Spatial Reasoning
We now give three examples of spatial reasoning tasks. Firstly, a model could pose a COP task where it challenges its competitor’s ability to reason about rotating a dice:
    class Dice:
        def __init__(self):
            self.faces = {
                "top": "1", "bottom": "6",
                "front": "2", "back": "5",
                "left": "3", "right": "4"
        def _rotate(self, plane, angle):
            """Some function to rotate the faces given a plane and angle"""
            self.faces = # update rules
        def get_face(face):
            return self.faces[face]
    d = Dice()
    d._rotate(’x’, 90)
    d._rotate(’y’, 270)
    print(d.get_face(’top’) + d.get_face(’left’))
This pseudocode was written by the authors. Secondly, consider task asking given a graph, can node A reach node B via a path?. Humans could solve this easily using spatial reasoning, and this is what we want to test. The challenge can be converted to COP using path search methods:
    from collections import deque
    graph = {
        "A": ["B","C"],
        "B": ["D"],
        "C": ["E"],
        "D": ["F"],
        "E": [],
        "F": []
    def reachable(start, goal):
        q = deque([start])
        seen = set()
        while q:
            u = q.popleft()
            if u == goal: return True
            for v in graph[u]:
                if v not in seen:
                    seen.add(v)
                    q.append(v)
        return False
    print(reachable("A","F"))
Next, the task is this 5-node graph 3-colorable? is converted to COP as follows:
    import itertools
    edges = [(0,1),(1,2),(2,3),(3,4),(4,0)]
    colors = [0,1,2]
    for assignment in itertools.product(colors, repeat=5):
        if all(assignment[u] != assignment[v] for u,v in edges):
            print("3-colorable:", assignment)
            break
    else:
        print("Not 3-colorable")
Lastly, the task How many paths are there from $(0,0)$ to $(6,6)$ that move only right or up, but don’t pass through $(3,3)$? can be converted as follows:
    from itertools import product
    def count_paths(n):
        paths = 0
        # Represent a path as a sequence of ’R’ and ’U’
        for seq in product("RU", repeat=2*n):
            if seq.count("R") == n and seq.count("U") == n:
                x = y = 0
                ok = True
                for move in seq:
                    if move == "R": x += 1
                    else: y += 1
                    if (x, y) == (3,3):
                        ok = False
                        break
                if ok:
                    paths += 1
        return paths
    print(count_paths(6))
C.4 Games/ Puzzles
We now consider two examples of games (connect-4 and chess). It is straightforward to convert interesting challenges within these settings into COP tasks.
Firstly, consider Connect-4 on a $6\times 7$ grid with some pieces already in play. The question is Can player X force a win in one move?. This can be converted into a COP task as follows:
    def next_empty_row(board, col):
    # next empty row for a particular col
    pass
    def check_win(board,row,col,player):
    # brute force check horizontal, vertical, diagonal for 4 in a row
    pass
    board = # set up initial board position
    for col in range(7):
        row = next_empty_row(board,col)
        board[row][col]="X"
        if check_win(board,row,col,"X"):
            print("Winning move in column",col)
        board[row][col]=None
This pseudocode was written by the authors. Secondly, consider a chess-based task: Assuming optimal behavior according to the Syzygy tablebase, given a board state X, what is the best move?. The Syzygy Chess engdame tablebase contains exact game-theoretic results of every legal position with up to 7 pieces. This can be written in COP-style as follows:
    import chess, chess.syzygy
    tb = chess.syzyg.open_tablebases("path/to/syzygy")
    # some board state
    board = chess.Board("8/8/8/8/8/8/1k6/K5R1 w - - 0 1")
    best = max(board.legal_moves, key=lambda m: tb.probe_wdl(board.copy(stack=False).push(m)))
    print(best)
Whilst our experiments excluded the use of external libraries, in this instance using the chess library would afford a wide option of chess-based tasks to be converted to COP (best move, set of legal moves, etc.).
Appendix D Question Clustering